The water drops fall at regular intervals from a tap 5 m above the ground. The fifth drop is leaving the tap at an instant when the first drop touches the ground. How far above the ground is the third drop at that instant? (Take g = 10 m/s-1)   
 

Grade:12th pass

1 Answers

Arun
4 years ago
Height =h=5m
s=ut+1/2at²
5=0+1/2aT²-----(1)
Time taken by second drop will be T/2
∴ its distance from tap
h=0+1/2(aT²)/4       ------------(2)
divide equ (ii) by equ (1)
h/5=1/4
h=1.25m
∴ H-h=5-1.25=3.75m